OpenStudy (anonymous):

graph the equation and identify the y-concept y+5y=15 where do both points go?

OpenStudy (anonymous):

both points? you mean y intercept and x intercept? assuming you meant "x + 5y = 15", you find the y intercept by setting x=0 and solving for y: 0 + 5y = 15 y = 3 you find the x intercept by setting y=0 and solving for x. x + 0 = 15 x = 15 So the two points are (0,3) and (15,0).

OpenStudy (anonymous):

here's the graph with the points: http://bit.ly/puP2iY

OpenStudy (anonymous):

so what is the y-intercept?

OpenStudy (anonymous):

it's the point where the graph crosses the y axis.